# The Data Acquisition System of the LZ Dark Matter Detector: FADR

### v2

@@ 1 Introduction
Interactions in the OD are observed with 120 8-inch diameter photomultiplier tubes (PMTs), mounted inside the water tank on stainless steel ladders. Events in the OD produce S1-like signals.
@@
RAID Reduntant Array of Independent Disks
@@ 2 The Signal Chain
Table 3: The response to SPHE signals from the TPC, Top and Bottom Skin, Dome, and OD PMTs, processed with the high-gain channels of the amplifier, at the input of the digitizer. The values reported here are average values and their standard deviations of all SPHE PMT signals obtained for each detector volume during LED calibrations. The TPC, Bottom Skin, Dome, and OD PMTs are operating at a gain of 2×$\times 10^{6}$ but use different values for the load resistor in the PMT base. The Top Skin PMTs are operating at a gain of 1×$\times 10^{6}$.
@@ 3 Digital Hardware
Each DDC-32 has two 14-bit, 100-MHz Analog-to-Digital Converter (DAC) ”spy” outputs, allowing for diagnostics and monitoring. Individual incoming channels or their digital sum can be send to the spy outputs. Digital-filter outputs, individual or summed, or any other signal internal to the FPGA can also be viewed using the spy outputs. This feature is very useful, especially during the development and deployment stages.
@@ 4.1 Digitization
Figure 10: The memory organization of the POD waveform storage in the FPGA for a single buffer (out of two) of a single channel. The separation of POD header and POD samples improves the performance and ease of extracting information from memory.
@@ 5.2 Data Sparsification
After an event is captured, a post-event holdoff is applied. During the holdoff period no triggers initiate event capture. The post-event holdoff is applied to reduce the probability of retriggering on the tails of a large S2 pulse. The holdoff period used during SR1 was 2 ms. The event selection parameters used during Science Run 1 (SR1) are listed in Table 7. Figure 15 shows an example of the S2 trigger operation.
@@ 8.2 DAQScope
DAQScope was developed to allow us to monitor individual signals connected to the DDC-32s without the need to move/unplug/replug any signal cables. The key elements of DAQScope are shown in Fig. 18. The operator can look at any one or a sum of selected PMT channels on a regular oscilloscope, connected to one of the spy outputs of the DDC-32s. The multiplexing/summing and digital-to-analog conversion is done at the full 100-MHz sample rate with no zero-suppression. The video output of the Tektronix scope is connected to a Video to IP Encoder [23]). This allows the individual signals to be monitored remotely, using VLC [24]. In the control room, an IP to Video Decoder [25] is used to display the scope image continuously on a large display. The DAQScope video stream is recorded, and the recordings are preserved for a limited time to allow delayed diagnostic of unusual detector behavior. The settings of DAQScope (e.g. the channels being displayed) are controlled via the Run Control interface.
Figure 18: Key elements of DAQScope. Two spy outputs are displayed on an oscilloscope in one of the electronics racks. The graphics output of the oscilloscope is connected to a Video to IP encoder. This allows the screen of the oscilloscope to be monitored by local and remote users. An IP to Video decoder is used to provide a permanent display of the image on the oscilloscope on a large display in the control room.
@@ 9.2 DDC-32 Linearity
The integral and differential nonlinearity of each DDC-32 channel was measured using a low distortion sine wave generator. For these measurements, we used the DC1858A from Analog Devices [26] which is suitable for testing ADCs with resolutions up to 18 bits. The information collected during these measurements is preserved and can be used to correct the measured data. The data collected for all 1632 DDC-32 channel are shown in Fig. 20. The integral nonlinearity is less than $\pm$3 ADCC. The differential nonlinearity varies between -0.5 ADCC and +1.0 ADCC.
@@ 9.3 System Noise
The electronic noise in the system is monitored in multiple ways. In this section two different approaches are discussed.
@@ 9.4 Data Transfer
Each channel of the DDC-32 has two circular buffers. Data can be written to one buffer while data from the other buffer are being transferred. For each buffer, the start and end times are captured in the data stream. The buffer start time is the time that a buffer is ready to receive data; the buffer end time is the time of the end of the event or when the first buffer of a particular digitizer is filled. Different digitizers can have different buffer start and end times. If the buffer stop time of event $n$ is larger than the buffer start time of event $n+1$ there is no dead time associated with switching buffers. If the buffer stop time of event $n$ is less than the buffer start time of event $n+1$ there is period when no data could be captured and dead time is incurred. This happens when one buffer is transferring data and the second buffer is waiting to transfer data. To offload one full buffer (18,412 samples) requires 340 $\mu$s. The maximum data transfer time when all 32 channels have a full buffer is 11 ms. The data from the 14 Data Extractors are transferred to 14 Data Collectors. Currently we can sustainably acquire data at 1,200 MB/s.
@@
To ensure data integrity we have implemented multiple levels of checksumming. First, every byte of data for each event is checksummed (CRC32) as it is read out from the Block RAMs of the DDC-32 FPGAs. The Data Extractors, as they split individual event data into jumbo frame-sized chunks with a maximum size of 8,800 bytes, checksum each UDP packet payload that is sent to the Data Collectors. The Data Collectors cross-check the payload checksums of each received UDP packet, and perform an event-level checksum cross-check once a given event is stitched back together. All the server machines in the data pipeline (Data Collectors, Event Builders, etc) use ECC-enabled memory modules and all file systems in the data chain use data and metadata checksumming. As ROOT sequence files are being generated on the Event Builders, their SHA256 checksums are calculated and stored in a data-tracking DB. These SHA256 checksums are recalculated and cross-checked at NERSC, before a sequence file is being marked as successfully transferred and permanently stored.
@@ 9.6 Event Selection
During WIMP Search, the OD waveforms are processed by an S1 filter with the following parameters:
@@ 9.7.2 POD Consistency
Two additional studies were carried out to verify the proper operation of the POD algorithm. We simulated the FPGA code to compare POD crossings observed in the raw waveforms to the PODs captured by the FPGA. To ensure that the POD algorithm did not miss any peaks, we compared the PODs captured by the FPGA with peaks identified in the raw waveforms by a peak finding algorithm that is part of the python package SciPy [27]. These studies used data files collected in the dual-capture mode discussed above with a POD threshold of 25 ADCC was used.
@@ 10 Summary
The design of the data acquisition system for LZ, FADR, was based on our experience with LUX and the required calibration rates and data volume for LZ. The RMS baseline noise of the entire signal processing chain for most PMTs is between 2.2 and 2.4 ADCC ($\approx$0.3 mV), ensuring a high efficiency for the detection of single photoelectrons with amplitudes of around 50 ADCC (6 mV). In order to reduce the impact of slow baseline variations, the PMT waveforms are processed with digital filters that perform automatic baseline subtraction and waveform integration before event selection. The design of FADR is optimized for data transfer. The data for a specific event are written to 14 Data Collectors. The overall performance of the system is limited by the Data Collector that sees the highest data volume. Currently we can sustainably acquire data at 1,200 MB/s.

### v3

@@ 1 Introduction
Interactions in the OD are observed with 120 8-inch diameter PMTs, mounted inside the water tank on stainless steel ladders. Events in the OD produce S1-like signals.
@@
RAID Redundant Array of Independent Disks
@@ 2 The Signal Chain
Table 3: The response to SPHE signals from the TPC, Top and Bottom Skin, Dome Skin, and OD PMTs, processed with the high-gain channels of the amplifier, at the input of the digitizer. The values reported here are average values and their standard deviations of all SPHE PMT signals obtained for each detector volume during LED calibrations. The TPC, Bottom Skin, Dome, and OD PMTs are operating at a gain of 2×$\times 10^{6}$ but use different values for the load resistor in the PMT base. The Top Skin PMTs are operating at a gain of 1×$\times 10^{6}$.
@@ 3 Digital Hardware
Each DDC-32 has two 14-bit, 100-MHz Analog-to-Digital Converter (DAC) ”spy” outputs, allowing for diagnostics and monitoring. Individual incoming channels or their digital sum can be sent to the spy outputs. Digital-filter outputs, individual or summed, or any other signal internal to the FPGA can also be viewed using the spy outputs. This feature is very useful, especially during the development and deployment stages.
@@ 4.1 Digitization
Figure 10: The memory organization of the POD waveform storage on the FPGA for a single buffer (out of two) of a single channel. The separation of POD header and POD samples improves the performance and ease of extracting information from memory.
@@ 5.2 Data Sparsification
After an event is captured, a post-event holdoff is applied. During the holdoff period no triggers initiate event capture. The post-event holdoff is applied to reduce the probability of retriggering on the tails of a large S2 pulse. The holdoff period used during LZ’s first Science Run (SR1) was 2 ms. The event selection parameters used during SR1 are listed in Table 7. Figure 15 shows an example of the S2 trigger operation.
@@ 8.2 DAQScope
DAQScope was developed to allow us to monitor individual signals connected to the DDC-32s without the need to move/unplug/replug any signal cables. The key elements of DAQScope are shown in Fig. 18. The operator can look at any one or a sum of selected PMT channels on a regular oscilloscope, connected to one of the spy outputs of the DDC-32s. The multiplexing/summing and digital-to-analog conversion is done at the full 100-MHz sample rate with no zero-suppression. The video output of the Tektronix scope is connected to a Video-to-IP Encoder [23]. This allows the individual signals to be monitored remotely, using VLC [24]. In the control room, an IP-to-Video Decoder [25] is used to display the scope image continuously on a large display. The DAQScope video stream is recorded, and the recordings are preserved for a limited time to allow delayed diagnostic of unusual detector behavior. The settings of DAQScope (e.g. the channels being displayed) are controlled via the Run Control interface.
Figure 18: Key elements of DAQScope. Two spy outputs are displayed on an oscilloscope in one of the electronics racks. The graphics output of the oscilloscope is connected to a Video-to-IP encoder. This allows the screen of the oscilloscope to be monitored by local and remote users. An IP-to-Video decoder is used to provide a permanent display of the image on the oscilloscope on a large display in the control room.
@@ 9.2 DDC-32 Linearity
The integral and differential nonlinearity of each DDC-32 channel was measured using a low distortion sine wave generator. For these measurements, we used the DC1858A from Analog Devices [26] which is suitable for testing ADCs with resolutions up to 18 bits. The information collected during these measurements is preserved and can be used to correct the measured data. The data collected for all 1632 DDC-32 channels are shown in Fig. 20. The integral nonlinearity is less than $\pm$3 ADCC. The differential nonlinearity varies between -0.5 ADCC and +1.0 ADCC.
@@ 9.3 System Noise
The electronic noise in the system is monitored in multiple ways. In this section, two different approaches are discussed.
@@ 9.4 Data Transfer
Each channel of the DDC-32 has two circular buffers. Data can be written to one buffer while data from the other buffer is being transferred. For each buffer, the start and end times are captured in the data stream. The buffer start time is the time that a buffer is ready to receive data; the buffer end time is the time of the end of the event or when the first buffer of a particular digitizer is filled. Different digitizers can have different buffer start and end times. If the buffer stop time of event $n$ is larger than the buffer start time of event $n+1$ there is no dead time associated with switching buffers. If the buffer stop time of event $n$ is less than the buffer start time of event $n+1$ there is period when no data could be captured and dead time is incurred. This happens when one buffer is transferring data and the second buffer is waiting to transfer data. To offload one full buffer (18,412 samples) requires 340 $\mu$s. The maximum data transfer time when all 32 channels have a full buffer is 11 ms. The data from the 14 Data Extractors are transferred to 14 Data Collectors. Currently we can sustainably acquire data at 1200 MB/s.
@@
To ensure data integrity we have implemented multiple levels of checksumming. First, every byte of data for each event is checksummed (CRC32) as it is read out from the Block RAMs of the DDC-32 FPGAs. The Data Extractors, as they split individual event data into jumbo frame-sized chunks with a maximum size of 8,800 bytes, checksum each UDP packet payload that is sent to the Data Collectors. The Data Collectors cross-check the payload checksums of each received UDP packet, and perform an event-level checksum cross-check once a given event is stitched back together. All the server machines in the data pipeline (Data Collectors, Event Builders, etc) use ECC-enabled memory modules and all file systems in the data chain use data and metadata checksumming. As ROOT sequence files are being generated on the Event Builders, their SHA256 checksums are calculated and stored in a data-tracking DB. These SHA256 checksums are recalculated and cross-checked at NERSC, before a sequence file is marked as successfully transferred and permanently stored.
@@ 9.6 Event Selection
During SR1, the OD waveforms are processed by an S1 filter with the following parameters:
@@ 9.7.2 POD Consistency
Two additional studies were carried out to verify the proper operation of the POD algorithm. We simulated the FPGA code to compare POD crossings observed in the raw waveforms to the PODs captured by the FPGA. To ensure that the POD algorithm did not miss any peaks, we compared the PODs captured by the FPGA with peaks identified in the raw waveforms by a peak finding algorithm that is part of the python package SciPy [27]. These studies used data files collected in the dual-capture mode discussed above with a POD threshold of 25 ADCC.
@@ 10 Summary
The design of the data acquisition system for LZ, FADR, was based on our experience with LUX and the required calibration rates and data volume for LZ. The RMS baseline noise of the entire signal processing chain for most PMTs is between 2.2 and 2.4 ADCC ($\approx$0.3 mV), ensuring a high efficiency for the detection of single photoelectrons with amplitudes of around 50 ADCC (6 mV). In order to reduce the impact of slow baseline variations, the PMT waveforms are processed with digital filters that perform automatic baseline subtraction and waveform integration before event selection. The design of FADR is optimized for data transfer. The data for a specific event are written to 14 Data Collectors. The overall performance of the system is limited by the Data Collector that sees the highest data volume. Currently we can sustainably acquire data at 1200 MB/s.